# On sufficient conditions for the transitivity of homeomorphisms

### v3

@@ 4. Proof of Theorem A
(a) Suppose that $A$ is dense in $X$ and $T$ has the barycenter property in $A$. Let $U$ and $V$ be nonempty open subsets of $X$. Since $A$ is dense in $X$, there are $p\in U$ and $q\in V$ such that $p,q\in A$; moreover, there exists $\varepsilon>0$ such that $B(p,\varepsilon)\subseteq U$ and $B(q,\varepsilon)\subseteq V$. By the barycenter property, there is a positive integer $N=N(\varepsilon,p,q)$ such that, for $n_{1}=1=n_{2}$, one can find an integer $m\in[0,N]$ and $x_{0}\in X$ satisfying $d\big{(}T^{i}(x_{0}),T^{i}(p)\big{)}<\varepsilon$ for every $-n_{1}\leqslant i\leqslant 0$ and $d\big{(}T^{i+m}(x_{0}),T^{i}(q)\big{)}<\varepsilon$ for every $0\leqslant i\leqslant n_{2}$. In particular, $d\big{(}x_{0},p\big{)}<\varepsilon$ and $d\big{(}T^{m}(x_{0}),q\big{)}<\varepsilon$. So, $T^{m}(U)\cap V\neq\emptyset$. Hence $T$ is topologically transitive.
@@
$d(T^{k}(x),a)\,<\,\frac{\varepsilon}{2}\quad\quad\text{and}\quad\quad d(T^{-\ell}(y),b)\,<\,\frac{\varepsilon}{2}.$ (4.1)
@@
$d(p,a)\,\,<\frac{\varepsilon}{2}\quad\quad\text{and}\quad\quad d(q,b)\,<\,\frac{\varepsilon}{2}.$ (4.2)
@@
$d(T^{-\ell}(y),q)$$\leqslant$ $d(T^{-\ell}(y),b)+d(q,b)\,<\,\frac{\varepsilon}{2}+\frac{\varepsilon}{2}\,=\,\varepsilon.$ (4.3)
@@
[…]
@@
$d\big{(}z,T(p)\big{)}\,<\,\varepsilon\quad\quad\text{and}\quad\quad d\big{(}T^{m}(z),T^{-1}(q)\big{)}\,<\,\varepsilon$ (4.4)
@@
$d(T^{k}(x),a)\,<\,\frac{\varepsilon}{2}\quad\quad\text{and}\quad\quad d(T^{-\ell}(y),b)\,<\,\frac{\varepsilon}{2}.$ (4.5)
@@
$d(p,a)\,\,<\frac{\varepsilon}{2}\quad\quad\text{and}\quad\quad d(q,b)\,<\,\frac{\varepsilon}{2}.$ (4.6)
@@
$d(T^{-\ell}(y),q)$$\leqslant$ $d(T^{-\ell}(y),b)+d(q,b)\,<\,\frac{\varepsilon}{2}+\frac{\varepsilon}{2}\,=\,\varepsilon.$ (4.7)
@@
$d(T^{n_{i}}(z_{i-1}),\,v_{i})\,<\,\varepsilon\quad\quad\text{and}\quad\quad d(T^{-m_{i}}(z_{i}),\,T(v_{i}))\,<\,\varepsilon.$ (4.8)
@@
Due to the inequalities (4.6), (4) and (4.8), the points $(x_{i})_{i=0}^{k+1+\ell+\sum_{i=1}^{m}(n_{i}+1+m_{i})}$ build an $\varepsilon-$chain linking $x$ to $y$. So, $T$ is chain-transitive. ∎
@@ 5. Proof of Theorem B
    $T$ has the weak $su-$intersecting property in $\operatorname{Per}(T)$.
@@ 6. Proof of Theorem C
$(1)$Rightarrow $(2)$Rightarrow $(3)$Rightarrow $(4)$Rightarrow $(5)$Rightarrow $\,(6)$Rightarrow $(7)$Rightarrow $(8)$Rightarrow $(9)$Rightarrow $(10)$
@@
$(4)\Rightarrow(5)$ Suppose that $f$ has the weak $su-$intersecting property in $\operatorname{Per}(f)$. According to [58, Lemma 2.2], $f$ has the barycenter property in $\operatorname{Per}(f)$. Since $f$ is Axiom $A$, one has $\Omega(f)\,=\,\overline{\operatorname{Per}(f)}$.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
By Lemma 6.2, $f$ has the barycenter property in $A$ for every $\emptyset\neq A\subseteq M$. Moreover, as $f$ is $\Omega-$stable, one has $\emptyset\neq\operatorname{Per}(f)=\operatorname{Per}_{h}(f)$. Therefore, $f$ has the barycenter property in $\operatorname{Per}_{h}(f)$, which, in this context, is equivalent to say that $f$ has the weak $su-$intersecting property in $\operatorname{Per}(f)$. ∎
@@
two-sided limit shadowing $\Rightarrow$ $su-$intersecting property in $M$. (6.1)
@@
specification propertyweak suintersecting property in $\text{specification property}\quad\Rightarrow\quad\text{weak $su-$intersectingproperty in $\operatorname{Per}_{h}(f)$}.$
@@ 7.2. Application 2
Since an Anosov diffeomorphism $f\colon M\to M$ is structurally stable (so the intersections of every stable manifold with every unstable manifold are transversal) and, by Theorem C, a topologically transitive Anosov diffeomorphism has the weak $su-$intersecting property in $\operatorname{Per}(f)$, we conclude that, under the transitivity assumption, every two periodic points of $f$ are homoclinically related. Consequently, by Theorem C we conclude that:
@@
If $f\in\mathcal{KS}$ and $f$ has the weak $su-$intersecting property in $\operatorname{Per}(f)$, then all the periodic points have the same index.
@@
Let $p,q$ be two periodic points of $f$. As $f$ has the weak $su-$intersecting property in $\operatorname{Per}(f)$, the sets $W^{s}_{f}(\mathcal{O}_{f}(p))\cap W^{u}_{f}(\mathcal{O}_{f}(q))$ and $W^{u}_{f}(\mathcal{O}_{f}(p))\cap W^{s}_{f}(\mathcal{O}_{f}(q))$ are nonempty. Taking into account that $f\in\mathcal{KS}$, we know that these intersections are transversal. Therefore, $p$ and $q$ are homoclinically related, and so they have the same index. ∎
@@
To simplify the notation, we will abbreviate the expression “$C^{1}$ interior of a set $S~{}\subseteq~{}{\rm Diff}^{1}(M)$” into $\operatorname{int}_{C^{1}}(S)$. Define
@@
$\mathcal{R}\,=\,\mathcal{KS}\cap\mathcal{I}.$ (7.1)
@@ 7.3. Application 3
Let $M$ be a smooth closed Riemannian manifold with dimension $\geqslant 2$ and $f\colon M\to M$ be a $C^{1}$ diffeomorphism. Recall that if $f$ has the weak $su-$intersecting property in $\operatorname{Per}(f)$, then $f$ has the barycenter property in $\operatorname{Per}(f)$; furthermore, if $\operatorname{Per}(f)=\operatorname{Per}_{h}(f)$, then $f$ has the weak $su-$intersecting property in $\operatorname{Per}(f)$ if and only if $f$ has the barycenter property in $\operatorname{Per}(f)$. Since in $\mathcal{KS}$ we have $\operatorname{Per}(f)=\operatorname{Per}_{h}(f)$, then generically $f$ has the weak $su-$intersecting property in $\operatorname{Per}(f)$ if and only if $f$ has the barycenter property in $\operatorname{Per}(f)$.
@@
[…]
@@
[…]
@@ 7.3.1. The $C^{1}$ interior of $\mathcal{WIP}_{h}$
The set $\operatorname{int}_{C^{1}}(\mathcal{WIP}_{h})$ coincides with the space of topologically transitive Anosov diffeomorphisms. Moreover, $\operatorname{int}_{C^{1}}(\mathcal{WIP}_{h})=\operatorname{int}_{C^{1}}(\mathcal{PCI})\cap\mathcal{CT}$.
@@
()($\subseteq$)( ⊆ ) Take $f\in\operatorname{int}_{C^{1}}(\mathcal{WIP}_{h})$. By Lemma 7.11, $f\in\operatorname{int}_{C^{1}}(\mathcal{PCI})$, and, by Proposition 7.4, $f$ is $\Omega-$stable. Since $f$ also has the weak $su-$intersecting property in $\operatorname{Per}(f)$, then $f$ is chain-transitive by Theorem C. So, $f\in\operatorname{int}_{C^{1}}(\mathcal{PCI})\cap\mathcal{CT}$.
()superset-of-or-equals($\supseteq$)( ⊇ ) Suppose that $f\in\operatorname{int}_{C^{1}}(\mathcal{PCI})\cap\mathcal{CT}$. By Corollary 7.6, $f$ is a topologically transitive Anosov diffeomorphism. As the space of topologically transitive Anosov diffeomorphisms is open, there exists a $C^{1}$ open neighborhood $\mathcal{U}$ of $f$ such that any $g\in\mathcal{U}$ is a topologically transitive Anosov diffeomorphism. By Theorem C, every $g\in\mathcal{U}$ has the weak $su-$intersecting property in $\operatorname{Per}(g)$. So, $f\in\operatorname{int}_{C^{1}}(\mathcal{WIP}_{h})$.
@@
$\operatorname{int}_{C^{1}}(\mathcal{GL})=\operatorname{int}_{C^{1}}(\mathcal{TLS})=\operatorname{int}_{C^{1}}(\mathcal{SPE})=\operatorname{int}_{C^{1}}(\mathcal{IP})=\operatorname{int}_{C^{1}}(\mathcal{IP}_{h})=\operatorname{int}_{C^{1}}(\mathcal{WIP}_{h}).$
@@
$\operatorname{int}_{C^{1}}(\mathcal{GL})$$\subseteq$ $\operatorname{int}_{C^{1}}(\mathcal{WIP}_{h}).$ (7.2)
Now, take $f\in\operatorname{int}_{C^{1}}(\mathcal{WIP}_{h})$ and $\mathcal{U}$ a $C^{1}$ open neighborhood of $f$ contained in $\mathcal{WIP}_{h}$. By Proposition 7.10, every $g\in\mathcal{U}$ is a topologically transitive Anosov diffeomorphism. In particular, every $g\in\mathcal{U}$ is $\Omega-$stable and has the weak $su-$intersecting property in $\operatorname{Per}(g)$. Thus, by Theorem C, every $g\in\mathcal{U}$ has the two-sided limit shadowing property, the specification property and the gluing orbit property. Hence,
@@
$\operatorname{int}_{C^{1}}(\mathcal{WIP}_{h})\,\subseteq\,\operatorname{int}_{C^{1}}(\mathcal{TLS})\,\cap\,\operatorname{int}_{C^{1}}(\mathcal{SPE})\,\cap\,\operatorname{int}_{C^{1}}(\mathcal{GL}).$ (7.3)

### v4

@@ 4. Proof of Theorem A
(a) Suppose that $A$ is dense in $X$ and $T$ has the barycenter property in $A$. Let $U$ and $V$ be nonempty open subsets of $X$. Since $A$ is dense in $X$, there are $p\in U$ and $q\in V$ such that $p,q\in A$; moreover, there exists $\varepsilon>0$ such that $B(p,\varepsilon)\subseteq U$ and $B(q,\varepsilon)\subseteq V$. By the barycenter property, there is a positive integer $N=N(\varepsilon,p,q)$ such that, for $n_{1}=1=n_{2}$, one can find an integer $m\in[0,N]$ and $x_{0}\in X$ satisfying $d\big{(}T^{i}(x_{0}),T^{i}(p)\big{)}<\varepsilon$ for every $n_{1}\leqslant i\leqslant 0$ and $d\big{(}T^{i+m}(x_{0}),T^{i}(q)\big{)}<\varepsilon$ for every $0\leqslant i\leqslant n_{2}$. In particular, $d\big{(}x_{0},p\big{)}<\varepsilon$ and $d\big{(}T^{m}(x_{0}),q\big{)}<\varepsilon$. So, $T^{m}(U)\cap V\neq\emptyset$. Hence $T$ is topologically transitive.
@@
(4.1) $d(T^{k}(x),a)\,<\,\frac{\varepsilon}{2}\quad\quad\text{and}\quad\quad d(T^{-\ell}(y),b)\,<\,\frac{\varepsilon}{2}.$
@@
(4.2) $d(p,a)\,\,<\frac{\varepsilon}{2}\quad\quad\text{and}\quad\quad d(q,b)\,<\,\frac{\varepsilon}{2}.$
@@
(4.3) $d(T^{-\ell}(y),q)$ $\leqslant$ $d(T^{-\ell}(y),b)+d(q,b)\,<\,\frac{\varepsilon}{2}+\frac{\varepsilon}{2}\,=\,\varepsilon.$
@@
$d\big{(}T^{i}(z),T^{i}(T(p))\big{)}$ $<$ $\varepsilon\quad\quad\forall\,-n_{1}\,\leqslant\,i\,\leqslant\,0$
$d\big{(}T^{i+m}(z),T^{i}(T^{-1}(q))\big{)}$ $<$ $\varepsilon\quad\quad\forall\,\,0\,\leqslant\,i\,\leqslant\,n_{2}.$
@@
(4.4) $d\big{(}z,T(p)\big{)}\,<\,\varepsilon\quad\quad\text{and}\quad\quad d\big{(}T^{m}(z),T^{-1}(q)\big{)}\,<\,\varepsilon$
@@
(4.5) $d(T^{k}(x),a)\,<\,\frac{\varepsilon}{2}\quad\quad\text{and}\quad\quad d(T^{-\ell}(y),b)\,<\,\frac{\varepsilon}{2}.$
@@
(4.6) $d(p,a)\,\,<\frac{\varepsilon}{2}\quad\quad\text{and}\quad\quad d(q,b)\,<\,\frac{\varepsilon}{2}.$
@@
(4.7) $d(T^{-\ell}(y),q)$ $\leqslant$ $d(T^{-\ell}(y),b)+d(q,b)\,<\,\frac{\varepsilon}{2}+\frac{\varepsilon}{2}\,=\,\varepsilon.$
@@
(4.8) $d(T^{n_{i}}(z_{i-1}),\,v_{i})\,<\,\varepsilon\quad\quad\text{and}\quad\quad d(T^{-m_{i}}(z_{i}),\,T(v_{i}))\,<\,\varepsilon.$
@@
Due to the inequalities (4.6), (4) and (4.8), the points
$(x_{i})_{i=0}^{k+1+\ell+\sum_{i=1}^{m}(n_{i}+1+m_{i})}$
build an $\varepsilon$-chain linking $x$ to $y$. So, $T$ is chain-transitive. ∎
@@ 5. Proof of Theorem B
    $T$ has the weak susu-intersecting property in $\operatorname{Per}(T)$.
@@ 6. Proof of Theorem C
$(1)$ $\Rightarrow$ $(2)$ $\Rightarrow$ $(3)$ $\Rightarrow$ $(4)$ $\Rightarrow$ $(5)$ $\Rightarrow$ (6)\,(6) $\Rightarrow$ $(7)$ $\Rightarrow$ $(8)$ $\Rightarrow$ $(9)$ $\Rightarrow$ $(10)$
@@
$(4)\Rightarrow(5)$ Suppose that $f$ has the weak susu-intersecting property in $\operatorname{Per}(f)$. According to [58, Lemma 2.2], $f$ has the barycenter property in $\operatorname{Per}(f)$. Since $f$ is Axiom $A$, one has $\Omega(f)\,=\,\overline{\operatorname{Per}(f)}$.
@@
$d\big{(}T^{j}(z),\,T^{j}(x)\big{)}$ $\leqslant$ $\varepsilon\quad\quad\forall\,0\leqslant j\leqslant n_{1}$
$d\big{(}T^{k+n_{1}+t_{1}}(z),\,T^{k}(y)\big{)}$ $\leqslant$ $\varepsilon\quad\quad\forall\,0\leqslant k\leqslant n_{2}.$
@@
$d\big{(}T^{j}(z_{\ell}),\,T^{j-\ell N}(p)\big{)}$ $=$ $d\big{(}T^{j}(z_{\ell}),\,T^{j}(x_{\ell})\big{)}\,\leqslant\,\varepsilon\quad\quad\forall\,0\leqslant j\leqslant n_{1,\ell}\,=\,\ell N$
$d\big{(}T^{k+\ell N+r_{\ell}}(z_{\ell}),\,T^{k}(q)\big{)}$ $=$ $d\big{(}T^{k+n_{1,\ell}+r_{\ell}}(z_{\ell}),\,T^{k}(y_{\ell})\big{)}\,\leqslant\,\varepsilon\quad\quad\forall\,0\leqslant k\leqslant n_{2,\ell}\,=\,\ell N.$
@@
$d\big{(}T^{k+\ell N+r_{\ell}-\ell N}(T^{\ell N}(z_{\ell})),\,T^{k}(q)\big{)}$ $=$ $d\big{(}T^{k+\ell N+r_{\ell}}(z_{\ell}),\,T^{k}(q)\big{)}\,\leqslant\,\varepsilon\quad\quad\forall\,0\leqslant k\leqslant\ell N.$
@@
$d\big{(}T^{k+r_{\ell}}(w_{\ell}),\,T^{k}(q)\big{)}$ $\leqslant$ $\varepsilon\quad\quad\forall\,0\leqslant k\leqslant\ell N.$
@@
By Lemma 6.2, $f$ has the barycenter property in $A$ for every $\emptyset\neq A\subseteq M$. Moreover, as $f$ is $\Omega$-stable, one has $\emptyset\neq\operatorname{Per}(f)=\operatorname{Per}_{h}(f)$. Therefore, $f$ has the barycenter property in $\operatorname{Per}_{h}(f)$, which, in this context, is equivalent to say that $f$ has the weak susu-intersecting property in $\operatorname{Per}(f)$. ∎
@@
(6.1) two-sided limit shadowing $\Rightarrow$ susu-intersecting property in $M$.
@@
$\text{specification property}\quad\Rightarrow\quad\text{weak $su-$intersecting property in $\operatorname{Per}_{h}(f)$}.$
@@ 7.2. Application 2
Since an Anosov diffeomorphism $f\colon M\to M$ is structurally stable (so the intersections of every stable manifold with every unstable manifold are transversal) and, by Theorem C, a topologically transitive Anosov diffeomorphism has the weak susu-intersecting property in $\operatorname{Per}(f)$, we conclude that, under the transitivity assumption, every two periodic points of $f$ are homoclinically related. Consequently, by Theorem C we conclude that:
@@
If $f\in\mathcal{KS}$ and $f$ has the weak susu-intersecting property in $\operatorname{Per}(f)$, then all the periodic points have the same index.
@@
Let $p,q$ be two periodic points of $f$. As $f$ has the weak susu-intersecting property in $\operatorname{Per}(f)$, the sets $W^{s}_{f}(\mathcal{O}_{f}(p))\cap W^{u}_{f}(\mathcal{O}_{f}(q))$ and $W^{u}_{f}(\mathcal{O}_{f}(p))\cap W^{s}_{f}(\mathcal{O}_{f}(q))$ are nonempty. Taking into account that $f\in\mathcal{KS}$, we know that these intersections are transversal. Therefore, $p$ and $q$ are homoclinically related, and so they have the same index. ∎
@@
To simplify the notation, we will abbreviate the expression
$C^{1}$ interior of a set $S~\subseteq~{\rm Diff}^{1}(M)$
into $\operatorname{int}_{C^{1}}(S)$.
Define
@@
(7.1) $\mathcal{R}\,=\,\mathcal{KS}\cap\mathcal{I}.$
@@ 7.3. Application 3
Let $M$ be a smooth closed Riemannian manifold with dimension $\geqslant 2$ and $f\colon M\to M$ be a $C^{1}$ diffeomorphism. Recall that if $f$ has the weak susu-intersecting property in $\operatorname{Per}(f)$, then $f$ has the barycenter property in $\operatorname{Per}(f)$; furthermore, if $\operatorname{Per}(f)=\operatorname{Per}_{h}(f)$, then $f$ has the weak susu-intersecting property in $\operatorname{Per}(f)$ if and only if $f$ has the barycenter property in $\operatorname{Per}(f)$. Since in $\mathcal{KS}$ we have $\operatorname{Per}(f)=\operatorname{Per}_{h}(f)$, then generically $f$ has the weak susu-intersecting property in $\operatorname{Per}(f)$ if and only if $f$ has the barycenter property in $\operatorname{Per}(f)$.
@@
$\mathcal{CT}$ $=$ $\big{\{}f\in{\rm Diff}^{1}(M)\colon\,\text{$f$ is chain-transitive}\big{\}}$
@@
$\mathcal{TLS}$ $=$ $\big{\{}f\in{\rm Diff}^{1}(M)\colon\,\text{$f$ has the two-sided limit shadowing property}\big{\}}$
@@ 7.3.1. The $C^{1}$ interior of $\mathcal{WIP}_{h}$
The set $\operatorname{int}_{C^{1}}(\mathcal{WIP}_{h})$ coincides with the space of topologically transitive Anosov diffeomorphisms.
Moreover, $\operatorname{int}_{C^{1}}(\mathcal{WIP}_{h})=\operatorname{int}_{C^{1}}(\mathcal{PCI})\cap\mathcal{CT}$.
@@
$(\subseteq)$ Take $f\in\operatorname{int}_{C^{1}}(\mathcal{WIP}_{h})$. By Lemma 7.11, $f\in\operatorname{int}_{C^{1}}(\mathcal{PCI})$, and, by Proposition 7.4, $f$ is $\Omega$-stable. Since $f$ also has the weak susu-intersecting property in $\operatorname{Per}(f)$, then $f$ is chain-transitive by Theorem C. So, $f\in\operatorname{int}_{C^{1}}(\mathcal{PCI})\cap\mathcal{CT}$.
$(\supseteq)$ Suppose that $f\in\operatorname{int}_{C^{1}}(\mathcal{PCI})\cap\mathcal{CT}$. By Corollary 7.6, $f$ is a topologically transitive Anosov diffeomorphism. As the space of topologically transitive Anosov diffeomorphisms is open, there exists a $C^{1}$ open neighborhood $\mathcal{U}$ of $f$ such that any $g\in\mathcal{U}$ is a topologically transitive Anosov diffeomorphism. By Theorem C, every $g\in\mathcal{U}$ has the weak susu-intersecting property in $\operatorname{Per}(g)$. So, $f\in\operatorname{int}_{C^{1}}(\mathcal{WIP}_{h})$.
@@
$\operatorname{int}_{C^{1}}(\mathcal{GL})=\operatorname{int}_{C^{1}}(\mathcal{TLS})=\operatorname{int}_{C^{1}}(\mathcal{SPE})=$
$\operatorname{int}_{C^{1}}(\mathcal{IP})=\operatorname{int}_{C^{1}}(\mathcal{IP}_{h})=\operatorname{int}_{C^{1}}(\mathcal{WIP}_{h}).$
@@
(7.2) $\operatorname{int}_{C^{1}}(\mathcal{GL})$ $\subseteq$ $\operatorname{int}_{C^{1}}(\mathcal{WIP}_{h}).$
Now, take $f\in\operatorname{int}_{C^{1}}(\mathcal{WIP}_{h})$ and $\mathcal{U}$ a $C^{1}$ open neighborhood of $f$ contained in $\mathcal{WIP}_{h}$. By Proposition 7.10, every $g\in\mathcal{U}$ is a topologically transitive Anosov diffeomorphism. In particular, every $g\in\mathcal{U}$ is $\Omega$-stable and has the weak susu-intersecting property in $\operatorname{Per}(g)$. Thus, by Theorem C, every $g\in\mathcal{U}$ has the two-sided limit shadowing property, the specification property and the gluing orbit property. Hence,
@@
(7.3) $\operatorname{int}_{C^{1}}(\mathcal{WIP}_{h})\,\subseteq\,\operatorname{int}_{C^{1}}(\mathcal{TLS})\,\cap\,\operatorname{int}_{C^{1}}(\mathcal{SPE})\,\cap\,\operatorname{int}_{C^{1}}(\mathcal{GL}).$